# ArcMemo: Abstract Reasoning Composition with Lifelong LLM Memory

### v2

@@ 3.5 Continual Concept Learning
A long-held goal in machine learning is to develop lifelong learning systems that continually self-improve without manual intervention. A memory system that can leverage the learning signal present at test time via continual updates is one approach to achieve this ambition.
@@ 4.1 Datasets
We evaluate our proposed framework on ARC-AGI-1 (Chollet, 2019), a benchmark explicitly designed to evaluate intelligence as “efficient acquisition of new skills” as opposed to “fixed possession/memorization of skills.” Each ARC puzzle encodes a transformation rule that maps input to output pixel grids. The objective of each puzzle is to infer its rule given several examples of input-output grid pairs, and produce the corresponding output grid for several input test cases. We find that the abstract domain of pixel grid transforms provides an interesting testbed to simulate frontier domains without requiring expert knowledge to evaluate trajectories–a confluence of desirable properties for evaluating a continual concept learning system.
ARC-AGI-1 contains a public validation split containing 400 puzzles with a difficulty distribution matching that of the private evaluation. Just as Akyürek et al. (2025) select an evaluation subset of 80 puzzles, our experiments evaluate against a uniform random subset of 100 puzzles from the public validation split to accommodate compute limitations.
Li et al. (2024) manually authored python solutions for 160 puzzles from the public train split (comparatively easier than public and private evaluation splits) to act as “seeds” to recombine for their synthetic dataset. We reuse these seeds but instead of initializing training data, we initialize our memory using these solutions (start with empty memory; call write operation with each seed puzzle and solution pair).
4.2 Settings
@@
To build on frontier models, we experiment primarily with OpenAI’s o4-mini. At the time of writing, o4-mini is second only to Grok 4, but o4-mini’s lower price puts it on the pareto frontier of cost and performance ARC-Prize (2025). For auxiliary tasks such as concept abstraction and non-reasoning selection, we use OpenAI’s GPT-4.1 to conserve token usage. Early experiments also evaluated the open-weight DeepSeek R1 which has the benefit of visible thinking traces, but its 8000 output token limit consistently yielded unfinished solutions in initial testing.
Li et al. (2024) demonstrated that the approaches of (1) induction: querying a model for a python program that takes an input grid and returns the transformed output grid and (2) transduction: querying a model to directly predict the output grids corresponding to the test case input grids are complementary. However since code solutions directly reflect the transformation steps and therefore provide more signal for reflection/abstraction, we use the induction method.
We take advantage of each puzzle’s reference examples and use them as test cases for generated programs. Using these test cases, we evaluate performance under 0, 1, and 2 retries with execution feedback. To mitigate sampling variance, we run each setting 3 independent times.
Our evaluation settings all use o4-mini (medium) and are as follows:
  • baseline: see sampling parameters in subsection C.2
  • cheatsheet: a re-implementation of Suzgun et al. (2025)’s Dynamic Cheatsheet as a baseline, that uses a frozen cheatsheet initialized against BARC seeds to match other settings.
  • ArcMemo-OE: our Open-Ended (OE) memory formulation
  • ArcMemo-PS: our Program Synthesis (PS) memory formulation
4.3 Scoring
We use ARC-AGI’s official scoring metric, oracle@$k$ (each test case is scored separately and if any of $k$ candidates pass, full credit is given for that test case). The ARC-AGI official evaluation has $k=2$, but to mitigate sampling variance, we sample 3 runs for each setting and report the average single run score, average oracle@2 score, and their standard deviation. We investigate several alternate scoring protocols (requiring a single program to solve all test cases, accumulating test case solve rates, and requiring references to also be solved) and include the results in Table 5.
@@ 5.1 Main Results
baseline (o4-mini) 46.33 (1.04) 55.17 (3.18)
cheatsheet 47.50 (2.78) 57.67 (2.52)
ArcMemo-OE (ours) 48.00 (1.00) 56.67 (1.53)
ArcMemo-PS (ours) 49.33 (0.29) 59.33 (0.29)
@@
External Scores
   BARC  [^2] Result is taken from Li et al. (2024) that fine-tunes 7B models and validates on the full validation split by selecting 2 predictions via majority vote from 20k samples for each puzzle. This is not directly comparable due the difficulty of normalizing compute between small open models and black-box API models. - 56.75
Official Leaderboard Scores  [^3] Official Leaderboard Scores are from ARC-Prize (2025) and also cannot be directly compared as the official leaderboard uses a private test set.
   Gemini 2.5 Pro (Thinking 16K) - 41.0
   Gemini 2.5 Pro (Thinking 32K) - 37.0
   Claude Opus 4 (Thinking 16K) - 35.7
   Claude Sonnet 4 (Thinking 16K) - 40.0
   o4-mini (medium) - 41.8
   o3 (medium) - 53.8
   o3-Pro (medium) - 57.0
Table 1: Main Results. ArcMemo-PS yields the best results with standard inference compute, and continues to scale with additional compute. Bracketed values represent standard deviation across 3 samples. o4-mini (medium) is used as the backbone for all tested methods, so we include “medium reasoning effort” external results from the official leaderboard.
As seen in Table 1, ArcMemo-PS achieves the best official score with standard compute  [^4] Matching the official evaluation scheme with 2 parallel attempts and 0 retries. and still benefits from additional scaling in both parallel samples and sequential retries. Table 4 shows complete details on all settings at all scales where ArcMemo-PS is the the only setting that consistently outperforms the baseline in all inference compute scales. Other memory formulations (cheatsheet, ArcMemo-OE) situationally improve over the baseline but still underperform it in some regimes.
We observe that ArcMemo-PS’s advantage is most prominent with lower compute regimes. This is consistent with the stated goal “memory for reasoning”: persisting previous experience aims to reduce redundancy from rediscovering ideas found in previous rollouts. The marginal impact of memory is reduced with more inference compute (attempts/tokens) as the model can rediscover ideas through exploration.
Comparing ArcMemo methods to our memory baseline (cheatsheet) shows that both methods are extremely competitive. Cheatsheet wins on oracle@2 at maximum depth, but is matched or exceeded by a ArcMemo setting in every other comparison. The quantitative improvement shown by ArcMemo is supported with qualitative analysis that finds ArcMemo memories to be more abstract and modular. This qualitative analysis (described more in subsection 5.2) observes that these abstract memories improve concept coverage– more target puzzle ideas are reflected in memory.
@@ 5.2 Selection Ablation and Qualitative Relevance Analysis
Table 2: Selection Ablation. Ablating the reasoning-based selection mechanism from ArcMemo-PS reduces overall performance, showing the selection mechanism is also useful for downstream performance.
We perform an ablation experiment removing the selection mechanism from ArcMemo-PS to examine its impact and to compare more directly to the selection-free Cheatsheet methodology. The generally better performance of the with-selection setting (as seen in Table 2) suggests that in addition to allowing memory selection to grow without overwhelming context in both abstraction phase/inference phase, it also helps downstream problem solving performance. Moreover, from the token usage graph in Figure 4, we see that concept selection’s performance improvement also comes at higher efficiency— the selection-ablated setting uses far more tokens. However there still exist certain scale regimes where the no-selection setting has a better score, suggesting that the selection mechanism has headroom for improvement (discussed more in section 6).
ArcMemo-PS (-selection) is most comparable to Dynamic Cheatsheet’s “cumulative” setting as all stored memory is included with each query. While both ultimately solve the same number of puzzles across all runs, we find that the settings differ in 10 puzzles (that is, each system solves 5 puzzles the other does not). We manually analyzed these puzzles and found that only 40% of new solves from Cheatsheet were related to memory elements seen in the cheatsheet while 100% of new solves from ArcMemo-PS-no-selection can be linked to concept memory contents. That is, we were unable to find relevant concepts in the cheatsheet to explain 60% of the new solves. We cannot definitively conclude that the ArcMemo was the causal factor in each of the new solves since LLM operations remain uninterpretable at scale, but taken at face value, this suggests that new solutions in the ArcMemo settings are more attributed to the ArcMemo itself as compared to simple sampling variance.
Manual inspection of a subset of reasoning-based selection results used for ArcMemo-PS, finds that while irrelevant concepts still appear within the selection, the key idea of the target is often included in the selection. The presence of a few distractor concepts appears to be manageable, as the reasoning model itself is capable of exploring with backtracking.
@@ 5.3 Continual Learning
Table 3: Continual Memory Learning. Comparing otherwise identical memory systems, we find that with sequential inference compute scale (at high retry depth), continually updating memory with new self-generated solutions leads to improved puzzle solving performance.
Results from our experiment comparing ArcMemo-OE with a version that updates during evaluation (every 10 problems; Table 3) show that our memory system’s performance improves with continual updates– a key requirement for meaningful lifelong learning. In particular, we find that the performance improvement emerges at later iterations in sequential inference scaling. We hypothesize this is because after more iterations and puzzle solving passes, new solutions are found, new memories are abstracted, and these new memories help solve other puzzles.
We briefly discuss additional items of token efficiency, concept specificity and embedding retrieval in Appendix B.
@@ 6 Conclusion
In this work, we introduce ArcMemo, a framework designed to support lightweight, lifelong learning on compositional reasoning tasks by emphasizing a higher level of abstraction and modularity. We explore two implementations of memory modules against ARC-AGI, a benchmark specifically designed to resist memorization and to evaluate fluid intelligence instead. Our main findings confirm the efficacy of our approach: ArcMemo-PS outscores comparable methods under official evaluation protocol and continues to benefit from additional inference compute scale. Moreover we observe that continual updates benefits memory augmentation over multiple attempts, and that selecting a subset of memory for each problem is a crucial component to enable a memory to continually grow without overwhelming LLM context. This paper’s work represents early attempts towards the main tenets of higher level abstraction and modularity. For future directions, we are most interested in designing a hierarchical consolidation mechanism (see Appendix A for discussion on limitations and future directions). While ArcMemo is a promising proof of concept, the abstract memory design space remains largely unexplored and we hope to encourage further investigation through the release of a concept-annotation dataset.
@@ Acknowledgements
Appendix A Limitations and Future Work
A.1 Limitations
Our work represents initial work into higher abstraction for text-based memory modules, and we hope encourages further exploration. For one, our study evaluates against a constrained “evaluation frontier”: we used a 100-puzzle subset of the ARC-AGI validation set, of which the non-memory baseline can already solve 73 with inference scale (maximum parallel attempts and sequential retries). This leaves only 27 puzzles where new solves are possible, and these tend to be highly challenging, limiting the absolute magnitude of gains we could observe. Second, introducing an updateable memory creates a potential sensitivity to the order of problems encountered during initialization or training. Although we freeze memory during evaluation, early updates can still influence its contents, and the final performance may depend on the sequence in which problems are processed. Understanding and controlling for this effect remains an open question.
A.2 Future Work
We see several promising directions for building on this work. One immediate opportunity is to extend evaluation to a broader and more challenging set of puzzles, including those unsolved by the baseline, to better quantify the upper bound of performance gains from memory augmentation and to test whether memory improves not just new solves but also pass@k consistency on borderline cases.
A second direction is to investigate order-robust update strategies and curriculum learning methods that reduce sensitivity to problem ordering while maintaining the benefits of inter-problem learning.
Another important avenue is hierarchical abstraction. While the existing PS formulation allows for new routines to be introduced as higher-level abstractions over existing routines or lower-level special cases of them, this additive approach cannot revise the concept structure as a whole. Moreover, our memory currently updates only when considering one solution at a time, whereas decisions about introducing new parent abstractions that encompass several existing ideas should require considering multiple experiences together. Future efforts could therefore focus on memory consolidation operations independent of individual puzzle-solving attempts.
Finally, our current memory format was designed through small-scale, cost-constrained experiments; richer optimization over representations, abstraction methods, and retrieval strategies could yield substantial improvements. To support this line of research, we release a dataset of frontier puzzles for o4-mini with hand-annotated concepts, along with helper puzzles demonstrating each concept at lower complexity. We hope this resource enables both hand-crafted and automated methods for finding optimal representations, exploring hierarchical abstraction, and developing memory consolidation operations independent of new inputs.
@@ A.2 Future Work
Table 4: Main Results. The right partition of the table containing aggregate scores show the inference scaling: going down represents sequential retry, going right represents adding parallel samples. ConceptMemPS is the only setting that outperforms the baseline consistently at every tested scale.
@@
Table 5: Strict Scoring. While the official evaluation scheme allows different test cases solved by different attempts to be ensembled, in contrast the strict scoring regime only makrs a puzzle as solved if a single attempt (generated program) solves all test cases.
Appendix B Further Discussion
@@
We find that our system’s token efficiency is similar to the baseline. We observe that memory-augmented runs tend to increase output token usage. While this seems to run counter to the intuition that retrieving rather than regenerating certain ideas would save tokens, we hypothesize that introducing memory leads the model to explore more hypotheses. In other words, selection inaccuracies (false positives) seem to be manageable by the reasoning models, but at the cost of increased token usage. Full results are plotted in Figure 4.
@@
As part of the development of our concept representations, we conducted small scale experiments using tiny (n=10) subsets of the validation split to investigate the level of detail needed for concepts to be useful. Manually writing maximum detail situation-suggestion style concepts and using iteratively LLM-summarized versions (for a total of 5 levels of specificity for each concept), we found the fairly expected result that higher levels of specificity correspond to better solve rates. Maximum specificity solved 4/10 with a generally decreasing pattern with lower specificity. While this result was entirely expected, the main goal of the experiment was to determine how much we could compromise on concept detail to improve the retrieval aspect. There is a tension between retrieval and puzzle solving in the sense that puzzle solving benefits from more detailed suggestions, but retrieval struggles to apply it to new situations when the concept’s relevant situation is so precisely defined.
@@
Early in testing, we investigated standard embedding-based retrieval approaches. Following the ArcMemo-OE setting, we used a VLM generated caption to query a vector database of ArcMemo-OE-style concept embeddings. While computationally cheap compared to autoregressive generation (especially so when leveraging long-form reasoning), we saw generally poor retrieval results throughout. Using OpenAI’s embeddings API along with an o3-mini puzzle solving backbone proved ineffective, lowering the score from 0.26 to 0.22– marking a 15% reduction. Qualitative analysis found that retrieved concepts were largely irrelevant to the puzzle at hand and seemed to over-index on lexical similarity. Moreover, a small set of broadly defined (non-specific) concepts were found to be grossly overrepresented.
Appendix C Implementation Details
C.1 PS Memory Format
Each PS format memory entry contains:
  • Title: a succinct label for the underlying idea.
  • Description: elaboration on behaviour and role.
  • Kind: whether this concept encodes a type/structure/routine.
  • Parameters: a list of fields that parametrize this concept.
C.2 Experiment Parameters
To build on frontier models, we experiment primarily with OpenAI’s o4-mini (max tokens=32000, reasoning_effort=medium). For auxiliary tasks such as concept abstraction and non-reasoning selection, we use OpenAI’s GPT-4.1 (temperature=0.3, max tokens=1000) to reduce token usage. We looked into evaluating the open-source DeepSeek R1 which also has transparent thinking process, but the output 8000 token limit consistently yielded unfinished solutions.
Appendix D Additional Related Work

### v3

@@ 3.5 Continual Concept Learning
A long-held goal in machine learning is to develop lifelong learning systems that continually self-improve without manual intervention. A memory system that can leverage the learning signal present at test-time via continual updates is one approach to achieve this ambition. Our memory write operations are lightweight queries that can ingest solution traces derived from both the system and external sources, making continual updates practical at scale.
@@ 4 Experiments
We evaluate our proposed framework on ARC-AGI-1 (Chollet, 2019), a benchmark explicitly designed to evaluate intelligence as “efficient acquisition of new skills” instead of “fixed possession/memorization of skills.” Each ARC puzzle encodes a transformation rule that maps input to output pixel grids. The objective of each puzzle is to infer its rule given several examples of input-output grid pairs, and produce the corresponding output grid for several input test cases. We find that the abstract domain of pixel grid transforms provides a meaningfully challenging testbed to simulate frontier domains without requiring expert knowledge to evaluate trajectories–a confluence of desirable properties for evaluating a continual concept learning system.
ARC-AGI-1 contains a public validation split containing 400 puzzles with a difficulty distribution matching that of the private evaluation. Following Akyürek et al. (2025), we evaluate a randomly selected 100-puzzle subset of the public val split. This makes repeated runs for more stable estimates feasible given cost and the sampling variance we observed. Li et al. (2024) manually authored Python solutions for 160 puzzles from the public train split to act as “seeds” to recombine into their synthetic dataset. We reuse these solutions to seed our memory rather than training data.
Models.
@@
To build on frontier models, we experiment primarily with OpenAI’s o4-mini. At the time of writing, o4-mini is second only to Grok 4, but o4-mini’s lower price puts it on the Pareto frontier of cost and performance ARC-Prize (2025). For auxiliary tasks such as concept abstraction and non-reasoning selection, we use OpenAI’s GPT-4.1 to conserve token usage. Early experiments also evaluated the open-weight DeepSeek R1, which has the benefit of visible thinking traces, but its $8000$ output token limit consistently yielded unfinished solutions in initial testing.
Evaluation.
While the official evaluation harness queries models to directly predict output grids for test cases, we instead use a program synthesis approach that queries for a transformation function to convert input to output grids. The code artifact provides more signal for reflection and also allows us to test proposed logic against reference pairs for feedback. We evaluate performance under 0, 1, and 2 retries with this execution feedback. We follow official ARC-AGI scoring (two attempts per puzzle), and account for sampling variance by averaging over extra runs (see details in Appendix B). The primary memory baseline we compare against is a re-implementation of Suzgun et al. (2025)’s DC-Cu (labeled “cheatsheet”) that uses a frozen memory to match other settings.
@@ 5.1 Main Results
qwen3-235b-a22b-instruct 11.00 (1.00) 16.67 (0.57)
deepseek r1 19.50 (7.05) 26.33 (4.02)
claude sonnet 4 (thinking 16k) 45.50 (1.00) 54.17 (0.58)
gemini 2.5 flash (thinking 16k) 9.67 (4.86) 13.83 (4.73)
o4-mini (medium) 46.33 (1.04) 55.17 (3.18)
cheatsheet (o4-mini (medium)) 47.50 (2.78) 57.67 (2.52)
ArcMemo-OE (ours, o4-mini (medium)) 48.00 (1.00) 56.67 (1.53)
ArcMemo-PS (ours, o4-mini (medium)) 49.33 (0.29) 59.33 (0.29)
@@
Table 1: Main Results. ArcMemo-PS yields the best results and scales with additional compute. Bracketed values represent standard deviation. See additional score details in Table 4.
As seen in Table 1, ArcMemo-PS achieves the best official score with standard compute  [^2] Matching the official evaluation scheme with 2 parallel attempts and 0 retries. and still benefits from additional scaling in both parallel samples and sequential retries. Table 4 shows complete details on all settings at all scales, where ArcMemo-PS is the only setting that consistently outperforms the baseline in all inference compute scales. Other memory formulations (Cheatsheet, ArcMemo-OE) situationally improve over the baseline but still underperform it in some regimes.
We observe that ArcMemo-PS’s advantage is most prominent with lower compute regimes. This is consistent with the stated goal “memory for reasoning”: persisting previous experience aims to reduce redundancy from rediscovering ideas found in previous rollouts. The marginal impact of memory is reduced with more inference compute as the model can rediscover ideas through exploration.
Compared to our memory baseline (Cheatsheet), ArcMemo methods are highly competitive, performing favorably in most regimes. The quantitative improvement shown by ArcMemo is supported by qualitative analysis that demonstrates ArcMemo memories are more modular. This qualitative analysis (described more in subsection 5.2) observes that these abstract memories improve concept coverage–more target puzzle ideas are reflected in memory. This difference may be explained by ArcMemo’s modular design being more conducive to maintaining a growing library.
@@ 5.2 Selection Ablation and Qualitative Relevance Analysis
Table 2: Selection Ablation. Ablating the reasoning-based selection mechanism from ArcMemo-PS reduces overall performance, showing selection is useful for downstream performance.
We perform an ablation experiment removing the selection mechanism from ArcMemo-PS to examine its impact and to compare more directly to the selection-free Cheatsheet methodology. The generally better performance of the with-selection setting (as seen in Table 2) suggests that, in addition to allowing memory to grow without overwhelming context in both the abstraction and inference phases, it also helps downstream problem-solving performance. Moreover, from the token usage plot in Figure 4, we see that concept selection’s performance improvement also comes at higher efficiency–the selection-ablated setting uses far more tokens. However, there still exist certain scale regimes where the no-selection setting has a better score. We attribute this to variance from imperfect selection and note that this indicates headroom for selection/retrieval.
ArcMemo-PS (-selection) is most comparable to Dynamic Cheatsheet’s “cumulative” setting (DC-Cu) as all stored memory is included with each query. While both ultimately solve the same number of puzzles across all runs, we find that the settings differ in 10 puzzles (that is, each system solves 5 puzzles the other does not). We manually analyzed these puzzles and found that only 40% of new solves in the cheatsheet setting were related to memory elements actually in the generated cheatsheet. That is, we were unable to find relevant notes in the cheatsheet to explain 60% of its new solves. In contrast, 100% of new solves from ArcMemo-PS (-selection) can be linked to concept memory contents. While we cannot definitively conclude ArcMemo changes directly induced each of the new solves (given the state of LLM interpretability), this seems to suggest new solves in the ArcMemo settings are potentially more attributable to the memory component compared to sampling variance.
Manual inspection of a subset of reasoning-based selection results used for ArcMemo-PS finds that while some irrelevant concepts still appear within the selection, the key idea of the target is often included in the selection. The presence of a few distractor concepts appears to be manageable, as the reasoning model itself is capable of exploring with backtracking.
@@ 5.3 Continual Learning
Table 3: Continual Memory Learning. Comparing otherwise identical memory systems, we find that with sequential inference compute scale (at high retry depth), continually updating memory with new self-generated solutions leads to improved puzzle-solving performance.
Results from our experiment comparing ArcMemo-OE with a version that updates during evaluation (every 10 problems; Table 3) show that our memory system’s performance improves with continual updates–a key requirement for meaningful lifelong learning. In particular, we find that the performance improvement emerges at later iterations in sequential inference scaling. We hypothesize this is because, after more iterations and puzzle-solving passes, new solutions are found, new memories are abstracted, and these new memories help solve other puzzles.
We further discuss token efficiency, concept specificity, and embedding retrieval in Appendix C, and limitations and future work discussion in Appendix D.
@@ 6 Conclusion
In this work, we introduce ArcMemo, a framework designed to support lightweight, lifelong learning on compositional reasoning tasks by emphasizing a higher level of abstraction and modularity. We explore two implementations of memory modules against ARC-AGI, a benchmark specifically designed to resist memorization and to evaluate fluid intelligence instead. Our main findings confirm the efficacy of our approach: ArcMemo-PS outscores comparable methods under the official evaluation protocol and continues to benefit from additional inference compute scale. Moreover, we observe that continual updates benefit memory augmentation over multiple attempts and that selecting a subset of memory for each problem is a crucial component to enable a memory to continually grow without overwhelming the LLM context. This paper’s work represents early attempts toward the main tenets of higher-level abstraction and modularity. Promising future directions include hierarchical designs and consolidation mechanisms that restructure memory. To encourage further investigation, we also release a concept-annotation dataset and configurable puzzle synthesis pipeline, providing resources for evaluating concept representations and advancing abstraction-based memory methods.
@@ Acknowledgements
Appendix A Implementation Details
A.1 PS Memory Format
Each PS format memory entry contains:
  • Title: a succinct label for the underlying idea.
  • Description: elaboration on behavior and role.
  • Kind: whether this concept encodes a type/structure/routine.
  • Parameters: a list of fields that parametrize this concept.
A.2 Experiment Parameters
To build on frontier models, we experiment primarily with OpenAI’s o4-mini (max tokens=32000, reasoning_effort=medium). For auxiliary tasks such as concept abstraction and non-reasoning selection, we use OpenAI’s GPT-4.1 (temperature=0.3, max tokens=1000) to reduce token usage. We looked into evaluating the open-source DeepSeek R1, which also has a transparent thinking process, but the output 8000 token limit consistently yielded unfinished solutions.
Appendix B Evaluation Details
We use ARC-AGI’s official scoring metric, which we term oracle@$k$ (each test case is scored separately, and if any of $k$ candidates pass, full credit is given for that test case). The ARC-AGI official evaluation has $k=2$, but to mitigate sampling variance, we sample 3 runs for each setting and report the average single run score, average oracle@2 score, and their standard deviation. We investigate several alternate scoring protocols (requiring a single program to solve all test cases, accumulating test case solve rates, and requiring references also to be solved) and include the results in Table 5.
Here is a precise definition of our scoring procedure for a single problem $P$. Let…
  • $X$ be the size $n$ set of attempts on problem $P$
  • $T$ be the set of test cases for problem $P$
  • $C$ represent a particular $k$-subset of $X$, in notation: $C\in[X]^{k}$
$score=\frac{1}{|[X]^{k}|}\sum\limits_{C\in[X]^{k}}z(C)$
where $z$ is the problem score by a $k$-attempt ensemble
$z(C)=\frac{1}{|T|}\sum\limits_{t_{i}\in T}\mathds{1}\{\exists c\in C\>\>\texttt{Verify}(c,t_{i})\}$
where
  • $c$ is a single program attempt in the k-attempt ensemble,
  • t is a test case,
  • and Verify returns true if the program successfully passes the test case.
@@ Appendix B Evaluation Details
Table 4: Full o4-mini Results. The right partition of the table containing aggregate scores shows the inference scaling: going down represents sequential retry, going right represents adding parallel samples. ArcMemo-PS is the only setting that outperforms the baseline consistently at every tested scale.
@@
Table 5: Strict Scoring. While the official evaluation scheme allows different test cases solved by different attempts to be ensembled, in contrast, the strict scoring regime only marks a puzzle as solved if a single attempt (generated program) solves all test cases.
Appendix C Further Discussion
@@
We find that our system’s token efficiency is similar to the baseline. We observe that memory-augmented runs tend to increase output token usage. While this seems to counter the intuition that retrieving rather than regenerating certain ideas would save tokens, we hypothesize that introducing memory leads the model to explore more hypotheses. In other words, selection inaccuracies (false positives) seem to be manageable by the reasoning models, but at the cost of increased token usage. Full results are plotted in Figure 4.
@@
As part of the development of our concept representations, we conducted small-scale experiments using tiny (n=10) subsets of the validation split to investigate the level of detail needed for concepts to be useful. Manually writing maximum detail situation-suggestion style concepts and using iteratively LLM-summarized versions (for a total of 5 levels of specificity for each concept), we found the reasonably expected result that higher levels of specificity correspond to better solve rates. Maximum specificity solved 4/10 with a generally decreasing pattern with lower specificity. While this result was entirely expected, the main goal of the experiment was to determine how much we could compromise on concept detail to improve the retrieval aspect. There is a tension between retrieval and puzzle solving in the sense that puzzle solving benefits from more detailed suggestions, but retrieval struggles to apply them to new situations when the concept’s relevant situation is so precisely defined.
@@
Early in testing, we investigated standard embedding-based retrieval approaches. Following the ArcMemo-OE setting, we used a VLM-generated caption to query a vector database of ArcMemo-OE-style concept embeddings. While computationally cheap compared to autoregressive generation (especially so when leveraging long-form reasoning), we saw generally poor retrieval results. Using OpenAI’s embeddings API and an o3-mini puzzle-solving backbone proved ineffective, lowering the score from 0.26 to 0.22, marking a 15% reduction. Qualitative analysis found that retrieved concepts were largely irrelevant to the puzzle and seemed to over-index on lexical similarity. Moreover, a small set of broadly defined (non-specific) concepts was found to be grossly overrepresented.
Figure 4: Token Efficiency Plot: Comparing various settings’ official score (oracle@2) on the public validation subset to the tokens used by the reasoning model for strictly puzzle solving.
Appendix D Limitations and Future Work
D.1 Limitations
Following Akyürek et al. (2025), our evaluation focuses on a 100-task subset of ARC-AGI-1, selected to make large-scale experimentation feasible under compute and cost constraints. This choice was further motivated by the substantial sampling variance we observed—identical prompts can yield noticeably different scores, requiring multiple runs to obtain stable estimates. As a result, our study concentrates on a relatively small frontier of puzzles where memory augmentation can yield new solves, limiting the absolute magnitude of observable gains.
In addition, ARC-AGI-2 was released mid-project as a harder successor benchmark, with state-of-the-art performance still below 30%. Re-running all baselines and experiments on ARC-AGI-2 was infeasible within scope due to cost and time constraints, but extending memory-based approaches to ARC-AGI-2 remains an important direction for future work.
D.2 Future Work
Several directions follow naturally from this work. One is to broaden evaluation to larger and more challenging puzzle sets, including those unsolved by the baseline, to test both new solve potential and consistency on borderline cases. Another is to explore order-robust update strategies and curriculum designs that reduce sensitivity to problem ordering while preserving inter-problem learning benefits. Beyond these extensions, a key avenue is hierarchical abstraction: moving from additive concept updates toward consolidation operations that restructure the memory across multiple experiences. Finally, our current design choices for representation and retrieval were made under cost constraints; richer optimization may yield substantial improvements.
To facilitate further research, we release a data resource: hand-annotated concepts for difficult puzzles, together with a configurable puzzle synthesis pipeline. This resource is intended to support isolated evaluation of concept representations and selection mechanisms, as well as development of methods for concept abstraction. We hope it provides a useful foundation for continued exploration of abstract reasoning memory.
Appendix E Additional Related Work